# 3D Motion Generation for Motion Language Models

### v1

@@ Abstract
This paper introduces VimoRAG, a novel video-based retrieval-augmented motion generation framework for motion large language models (LLMs). As motion LLMs face severe out-of-domain/out-of-vocabulary issues due to limited annotated data, VimoRAG leverages large-scale in-the-wild video databases to enhance 3D motion generation by retrieving relevant 2D human motion signals. While video-based motion RAG is nontrivial, we address two key bottlenecks: (1) developing an effective motion-centered video retrieval model that distinguishes human poses and actions, and (2) mitigating the issue of error propagation caused by suboptimal retrieval results. We design the Gemini Motion Video Retriever mechanism and the Motion-centric Dual-alignment DPO Trainer, enabling effective retrieval and generation processes. Experimental results show that VimoRAG significantly boosts the performance of motion LLMs constrained to text-only input. All the resources (https://github.com/WalkerMitty/VimoRAG) will be open soon.
@@ 1 Introduction
Generating diverse and realistic human motions from free-form text prompts has significant practical applications, including video gaming, robotic assistance, and virtual reality. Previous advancements, ranging from transformer-based VAEs t2m-gpt to recent diffusion-based generative models motion_diffuse , have led to an increasingly promising generative performance. With the emergence of LLMs, motion-language models (aka. motion LLMs) have been proposed motiongpt_aaai ; motiongpt_nips . These unified architectures not only understand various motions but also support motion generation. To achieve competitive capabilities, motion LLMs require training on extremely large-scale datasets to ensure sufficient capacity. Particularly for motion generation, a substantial amount of labeled data (i.e., text-motion pairs) is essential, without which, models face severe out-of-domain (OOD) and out-of-vocabulary (OOV) issues, making it difficult to generalize to the vast variety of dynamic human motions.
However, existing datasets of text-motion pairs are severely limited in scale, comprising only approximately 14k samples humanml3d , while the cost of large-scale annotation is prohibitively high. To address this issue, Zhang et al. introduce ReMoDiffuse remodiffuse , a retrieval-augmented generation (RAG) method that retrieves relevant supplementary supervision signals from 3D motion databases. While this method provides a promising direction to address the scarcity of labeled data, their effectiveness might still remain constrained by the size of existing 3D motion databases, i.e., totaling only 14k samples from datasets such as HumanML3D humanml3d .
To address this, this paper explores an innovative RAG paradigm: retrieving information from larger-scale in-the-wild videos to supplement abundant motion signals that can enhance motion generation. Although videos represent a 2D visual modality, intuitively, the 2D human motions depicted in videos inherently share similar characteristics with 3D human motions qiu2024vimo , which can be utilized to guide the learning process of motion LLMs. Most importantly, existing video data is highly accessible and virtually unlimited in scale, and in-the-wild videos capture diverse and unconstrained human motions, offering strong potential to address OOD/V challenges. To this end, we introduce a simple but effective framework, named VimoRAG (cf. Fig. 2). VimoRAG first retrieves a video from an unlabeled video database based on the input text, then inputs both the text and the retrieved videos into an LLM to generate motion tokens, which are finally decoded into a motion sequence using VQ-VAE vq-vae .
@@
We leverage a lightweight LLM Phi3-3.8B phi-3 following Maaz et al. videogpt+ to evaluate the performance of this framework. To evaluate its effectiveness, we conduct both cross-domain and in-domain experiments. To explore its potential, we conduct scaling experiments with varying retrieval corpus sizes. Specifically, in OOD scenarios, we conduct zero-shot experiments on the challenging IDEA400 motionx test set, where we verify that VimoRAG exhibits strong generalization capabilities. We evaluate in-domain performance on the representative HumanML3D benchmark and observe that VimoRAG consistently improves most of the metrics compared to existing motion LLMs that operate with text-only input, and further exhibit sustained performance gains as the retrieval corpus expands. In summary, our contributions are as follows:
@@
Figure 2: A high-level overview of our VimoRAG framework architecture.
@@ 2 Related Work
Motion generation motion_survey aims to generate human-like 3D motion based on a given prompt, such as text momask ; remodiffuse ; omg ; motionclip ; plan_posture ; fan2024textual ; oohmg . Several motion specialists t2m-gpt ; mdm ; mld achieve strong performance through carefully designed in-house training. In recent years, the emergence of LLMs openagi has demonstrated unprecedented levels of intelligence, driving the evolution from specialists to generalists. In the motion domain, numerous motion language models motiongpt_aaai ; motiongpt_nips ; motionllm ; motiongpt2 have been proposed. To achieve robust motion manipulation capabilities, these motion LLMs require fine-tuning on large annotated datasets. However, annotated datasets remain limited due to the high cost of annotation. To address this, Zhang et al. remodiffuse proposes a RAG-based solution, which retrieves related motion signals from a labeled motion database to enhance generation quality. In contrast, we find that video databases are significantly larger and more accessible. Thus, we consider a human motion-centric video retrieval approach to strengthen 3D motion generation, where the motion-persistent 2D features from videos can be aggressively transferred to the 3D motion process.
Notably, several motion LLM studies motionbank ; lmm ; motionllm have also explored human-related video tasks. Inspired by MotionBank motionbank , we construct our video corpus from multiple action-centric datasets. Unlike previous works that focus on building high-quality video collections, this work instead centers on validating the potential and robustness of VimoRAG framework when retrieving from a wild video corpus, and on addressing the potential inconsistency between video input and the generated motion.
@@ 3.1 Preliminaries
In theory, it is possible to retrieve the most semantically relevant videos from all available short videos on the internet to overcome the challenges of open-vocabulary text descriptions in the industry. However, in academia, in order to efficiently verify the feasibility of our method and to advance this research direction, we gather and filter a vast human-centric video database (HcVD) consisting of 425,988 videos, sourced from motionx ; aslan ; hmdb51 ; kinetics-400 ; penn_action ; ucf-101 ; nturgb .
To train a better retrieval model on human-centric videos, we leverage the widely used Qwen2-VL qwen2-vl model to synthesize textual descriptions following Zhao et al humanomni . We clarify that the synthetic captions are used solely to train the retriever, and are not involved in the motion RAG pipeline. VimoRAG is a ready-to-use framework that requires neither large-scale annotated videos nor text-to-text retrieval. Additionally, we enhance the dataset quality by employing AlphaPose alphapose to filter out videos without human detection. More details are presented in the Appendix C.
@@ 3.2 Gemini Motion Video Retriever
Figure 3 illustrates the overall architecture of our method. The model builds upon the CLIP clip framework, where the text and video branches are encoded separately, and their similarity is computed for subsequent ranking.
Our model consists of two independent retrievers—object-level and action-level—whose outputs are fused by a lightweight router to produce the final similarity score. Each retriever encodes both text and video with a specific focus. The object-level retriever captures visual entities (objects) and their textual arguments, while the action-level retriever targets motion and predicate-level semantics.
@@
Specifically, for the action-level retriever, we carefully design an action encoder $\theta_{\mathcal{A}}$ and predicate semantic extractor $\theta_{\mathcal{P}}$. In the action encoder, we first extract the 2D human keypoints from the $m$ frames of an input video $v$. Then we encode each frame of keypoints to the feature space separately. After this, we add learnable position embeddings frame by frame and feed them into a temporal encoder. In order to avoid information loss in the sequential encoding process, the residual operation is adopted here. To obtain a good initial weight, we adopt the pretrained AlphaPose alphapose and the pretrained MotionBERT motionbert as the keypoints detector and encoder accordingly. The temporal encoder is implemented by a transformer block transformer and is initialized randomly. $\theta_{\mathcal{P}}$ is initialized from the text encoder in InternVideo internvideo .
We train the action-level retriever using the contrastive learning loss clip4clip $\mathcal{L}_{action}=\mathcal{L}_{p2a}+\mathcal{L}_{a2p}$ ($\mathcal{L}_{a2p}$ is provided in the Appendix D due to space limitations):
@@
where $\mathbf{p}$ and $\mathbf{a}$ are embedding vectors encoded by $\theta_{\mathcal{P}}$ and $\theta_{\mathcal{A}}$ for the input value $t$ and $v$, respectively. $B$ denotes the number of text-video pairs in a training batch. $s(,)$ denotes the cosine similarity. For the object-level retriever, we directly adopt InternVideo internvideo , one of the most widely used VFMs, owing to its extensive common knowledge. The argument semantic extractor $\theta_{\mathcal{G}}$ and object encoder $\theta_{\mathcal{O}}$ are initialized using the text encoder and video encoder of InternVideo. During the fine-tuning stage, we utilize a loss function $\mathcal{L}_{object}$ similar to $\mathcal{L}_{action}$. Let g and o denote the embedding vectors encoded by $\theta_{\mathcal{G}}$ and $\theta_{\mathcal{O}}$ for the input value $t$ and $v$, respectively. We just replace p with g and replace a with o in $\mathcal{L}_{action}$, and then we can get the symmetric loss function $\mathcal{L}_{object}$. It is worth emphasizing that the two semantic extractors do not explicitly extract the semantics of predicates and arguments. We hypothesize that, through contrastive learning with different video features, each semantic extractor implicitly captures its respective focus—predicates and arguments.
@@
Figure 4: The McDPO training strategy. Given a text $t$ and a retrieved video $v$, we first perform visual demonstration-enhanced instruction tuning to establish a base reference model $\pi_{ref}$. Using $\pi_{ref}$ and the proposed motion-centric dual-alignment reward model, we construct a preference training set, and then apply DPO training on this dataset.
@@ 3.3 Motion-centric Dual-alignment DPO Trainer
Given an input text $x=\{x_{1},x_{2},...x_{n_{x}}\}$ of $n_{x}$ tokens, a system prompt $\mathcal{S}$ of the LLM, an instruction template $\mathcal{T}$, and a retrieved video $v=\{v_{1},v_{2},...,v_{n_{v}}\}$ of $n_{v}$ frames, we first embed $v$ into $k$ segment-wise embeddings $E^{v}=\{E^{v}_{1},E^{v}_{2},...,E^{v}_{k}\}$ following Maaz et al. videogpt+ We then concatenate these elements to obtain the final input embeddings $E^{f}=[emb(\mathcal{P}),\mathcal{T}(emb(x),E^{v})]$ where $emb(\cdot)$ denotes the embedding layers of the LLM. For the target motion, we leverage the widely used VQ-VAE vq-vae to encode the continuous motion sequence into discrete motion tokens $y=\{y_{1},y_{2},...,y_{n_{y}}\}$ following MotionGPT motiongpt_aaai . Inspired by them, we fine-tune the LLM using the following instruction-tuning format in Stage 1: System Prompt $\mathcal{P}$: You are a helpful AI assistant.
@@
MoMask momask 5.982±.089 0.110±.003 0.195±.006 0.266±.006 5.625±.023 7.558±.119
T2M-GPT t2m-gpt 5.359±.078 0.108±.006 0.186±.005 0.255±.006 5.773±.037 7.648±.100
MDM mdm 5.907±.107 0.113±.004 0.200±.004 0.278±.004 6.013±.020 8.131±.080
MotionDiffuse motion_diffuse 5.485±.038 0.110±.002 0.194±.002 0.266±.003 6.038±.005 6.884±.095
MLD mld 5.410±.085 0.114±.003 0.200±.005 0.270±.004 6.005±.029 7.558±.086
MotionGPT motiongpt_nips 6.202±.186 0.087±.005 0.151±.007 0.209±.008 6.640±.025 7.684±.111
ReMoDiffuse remodiffuse 9.639±.069 0.110±.004 0.188±.006 0.256±.005 5.465±.015 7.540±.120
@@
To construct such a training set, we first use $\pi_{ref}$ to randomly sample $\kappa$ times to generate a motions’ candidate set. In contrast to recent works sheng2024explorin ; pappa2024modipo , which rely on human or AI-based proxies as reward models, we devise a more efficient dual-alignment reward model. This model outputs a reward score for a sampled motion sequence $\hat{y}_{i}$ as follows:
@@
where $\ell(\cdot)$ and $d(\cdot)$ denote the distribution distance and Euclidean distance computed based on the features of the two inputs, respectively. $w_{\ell}$ and $w_{d}$ are hyper-parameters to respectively control the degree of alignment within the motion modality and between the text-motion modalities. The reward model encourages the preferred motions to be closer to the reference motions in the feature space and more semantically aligned with the paired text in the semantic space. By leveraging the reward model, we collect the chosen motions $y_{w}$ and the rejected motions $y_{l}$ from the existing motions’ candidate set, thereby constructing a DPO dataset $\mathcal{D}_{dpo}=\{(x,v,y^{w},y^{l})\}$. Finally, we adopt the following DPO training objective dpo ; llava-hound . Here $\pi_{\theta}$ and $\sigma$ denote the policy model and the logistic function, respectively. $\gamma$ here denotes the weighting coefficient.
@@
We employ phi-3-mini phi-3 as the backbone LLM and utilize the LoRA lora tuning method in both training stages. In Stage 1, we additionally tune the visual adapter while freezing the remaining modules. In Stage 2, all modules except for the LLM are kept frozen. Due to space constraints, further training configurations are detailed in the Appendix D.
Table 2: Results on HumanML3D test set. “*” denotes results from original papers; others are reproduced using official code. VimoRAG achieves the best FID and competitive performance across metrics among existing motion LLMs. This framework significantly improves five metrics (highlighted in red) over MotionGPT motiongpt_aaai (Phi3-3.8B), demonstrating the substantial advantage of incorporating video priors. The complete results with confidence intervals are shown in Table 7.
@@
MoMask momask 0.048 0.519 0.715 0.809 2.955 9.632
T2M-GPT t2m-gpt 0.112 0.489 0.679 0.774 3.125 9.691
MDM mdm 0.454 0.419 0.606 0.712 3.636 9.449
MotionDiffuse motion_diffuse 0.672 0.492 0.685 0.784 3.085 9.499
MLD mld 0.425 0.468 0.656 0.759 3.266 9.698
ReMoDiffuse remodiffuse 0.125 0.493 0.676 0.775 3.047 9.211
LMM* lmm 0.040 0.525 0.719 0.811 2.943 9.814
MotionLab* motionlab 0.167 0.810 2.912 9.593
MotionLCM* motionlcm 0.304 0.502 0.698 0.798 3.012 9.607
MotionCLR* motionclr 0.269 0.544 0.732 0.831 2.806
MotionGPT* motiongpt_nips 0.232 0.492 0.681 0.778 3.096 9.528
BiPO* bipo 0.030 0.523 0.714 0.809 2.880 9.556
StableMoFusion* stable_mofusion 0.098 0.553 0.748 0.841 9.748
MoGenTS* mogents 0.033 0.529 0.719 0.812 2.867 9.570
LAMP* lamp 0.032 0.557 0.751 0.843 2.759 9.571
@@
MotionGPT-2* motiongpt2 Llama3-8B 0.191 0.496 0.691 0.782 3.080 9.860
MotionLLM* motionagent GPT4+Gemma-2B 0.230 0.515 0.801 2.967 9.908
Wang et al.largemotionmodel Llama2-13B 0.166 0.519 0.803 2.964
ScaMo* scamo codesize 512-3B 0.617 0.443 0.627 0.734 3.340 9.217
AvatarGPT* avatargpt Llama-13B 0.567 0.389 0.539 0.623 9.489
MotionGPT* motiongpt_aaai Llama-13B 0.567 3.775 9.006
MotionGPT motiongpt_aaai Phi3-3.8B 0.501 0.396 0.575 0.673 3.724 9.475

### v2

@@ Abstract
This paper introduces VimoRAG, a novel video-based retrieval-augmented motion generation framework for motion large language models (LLMs). As motion LLMs face severe out-of-domain/out-of-vocabulary issues due to limited annotated data, VimoRAG leverages large-scale in-the-wild video databases to enhance 3D motion generation by retrieving relevant 2D human motion signals. While video-based motion RAG is nontrivial, we address two key bottlenecks: (1) developing an effective motion-centered video retrieval model that distinguishes human poses and actions, and (2) mitigating the issue of error propagation caused by suboptimal retrieval results. We design the Gemini Motion Video Retriever mechanism and the Motion-centric Dual-alignment DPO Trainer, enabling effective retrieval and generation processes. Experimental results show that VimoRAG significantly boosts the performance of motion LLMs constrained to text-only input. All the resources (https://walkermitty.github.io/VimoRAG/) are available.
@@ 1 Introduction
Generating diverse and realistic human motions from free-form text prompts has significant practical applications, including video gaming, robotic assistance, and virtual reality. Previous advancements, ranging from transformer-based VAEs [1] to recent diffusion-based generative models [2], have led to an increasingly promising generative performance. With the emergence of LLMs, motion-language models (aka. motion LLMs) have been proposed [3, 4]. These unified architectures not only understand various motions but also support motion generation. To achieve competitive capabilities, motion LLMs require training on extremely large-scale datasets to ensure sufficient capacity. Particularly for motion generation, a substantial amount of labeled data (i.e., text-motion pairs) is essential, without which, models face severe out-of-domain (OOD) and out-of-vocabulary (OOV) issues, making it difficult to generalize to the vast variety of dynamic human motions.
However, existing datasets of text-motion pairs are severely limited in scale, comprising only approximately 14k samples [5], while the cost of large-scale annotation is prohibitively high. To address this issue, Zhang et al. introduce ReMoDiffuse [6], a retrieval-augmented generation (RAG) method that retrieves relevant supplementary supervision signals from 3D motion databases. While this method provides a promising direction to address the scarcity of labeled data, their effectiveness might still remain constrained by the size of existing 3D motion databases, i.e., totaling only 14k samples from datasets such as HumanML3D [5].
To address this, this paper explores an innovative RAG paradigm: retrieving information from larger-scale in-the-wild videos to supplement abundant motion signals that can enhance motion generation. Although videos represent a 2D visual modality, intuitively, the 2D human motions depicted in videos inherently share similar characteristics with 3D human motions [7], which can be utilized to guide the learning process of motion LLMs. Most importantly, existing video data is highly accessible and virtually unlimited in scale, and in-the-wild videos capture diverse and unconstrained human motions, offering strong potential to address OOD/V challenges. To this end, we introduce a simple but effective framework, named VimoRAG (cf. Fig. 2). VimoRAG first retrieves a video from an unlabeled video database based on the input text, then inputs both the text and the retrieved videos into an LLM to generate motion tokens, which are finally decoded into a motion sequence using VQ-VAE [8].
@@
We leverage a lightweight LLM Phi3-3.8B [9] following Maaz et al. [10] to evaluate the performance of this framework. To evaluate its effectiveness, we conduct both cross-domain and in-domain experiments. To explore its potential, we conduct scaling experiments with varying retrieval corpus sizes. Specifically, in OOD scenarios, we conduct zero-shot experiments on the challenging IDEA400 [11] test set, where we verify that VimoRAG exhibits strong generalization capabilities. We evaluate in-domain performance on the representative HumanML3D benchmark and observe that VimoRAG consistently improves most of the metrics compared to existing motion LLMs that operate with text-only input, and further exhibit sustained performance gains as the retrieval corpus expands. In summary, our contributions are as follows:
@@
Figure 2: Overview of the VimoRAG pipeline: (1) text-to-video retrieval via Gemini-MVR, and (2) video-augmented motion generation guided by text and retrieved video. Gemini-MVR (Sec. 3.2) is designed to improve cross-modal human-centric video retrieval, while the McDPO training strategy (Sec. 3.3) mitigates error propagation caused by noisy retrievals.
@@ 2 Related Work
Motion generation [12] has long been a hot research topic in the related community, aiming to generate human-like 3D motion based on a given prompt, such as text, action, or incomplete motion. Text-to-motion is among the most significant task settings and has attracted substantial research attention [13, 6, 14, 15, 16, 17, 18]. For instance, T2M-GPT [1] explores a generative framework utilizing VQ-VAE [8] and Transformer [19] for motion generation. MDM [20] introduces a diffusion-based [21] generative model trained across multiple motion tasks. MLD [22] enhances the latent diffusion model to produce motions conditioned on various inputs. These motion generation specialists, following in-house training, deliver high performance in motion generation.
In recent years, the emergence of LLMs [23] has demonstrated unprecedented levels of intelligence, driving the evolution from specialists to generalists. In the motion domain, motion language models [3, 4, 24, 25] have been proposed, where motion-aware encoders are connected to a central LLM, leading to motion generalists (also named motion LLMs) capable of perceiving various motions. Further, motion generators (e.g., diffusion or VQ-VAE models) are integrated to achieve unified motion LLMs for both comprehension and generation [3]. To achieve robust motion manipulation capabilities, these motion LLMs require fine-tuning on large annotated datasets. However, compared to comprehension tasks, motion generation is more reliant on data, but motion annotation datasets are often quite limited due to the high cost of annotation.
In a recent study, Remodiffuse [6] introduces a motion generation method based on the retrieval-augmented generation (RAG) paradigm. It performs text-to-text retrieval from a labeled 3D motion database to fetch related motion signals and enhance generation quality. However, as previously mentioned, existing motion databases are typically limited in scale. In contrast, large-scale video databases are more accessible and diverse. Motivated by this, we explore a human motion-centric video retrieval framework to support 3D motion generation, where motion-consistent 2D features extracted from videos are effectively transferred to guide the 3D motion synthesis. Compared to Remodiffuse, our approach introduces two key innovations. First, we leverage cross-modal text-to-video retrieval to eliminate the reliance on motion databases that require manually written textual descriptions. Second, we are the first to identify the issue of error propagation in motion-RAG frameworks, and propose a novel algorithm, McDPO, to address it.
Notably, several motion LLM studies [26, 27, 24] have also explored human-related video tasks. Inspired by MotionBank, we construct our video corpus from multiple action-centric datasets. Unlike previous works that focus on building high-quality video collections, this work instead centers on validating the potential and robustness of VimoRAG framework when retrieving from a wild video corpus, and on addressing the potential inconsistency between video input and the generated motion.
@@ 3.1 Preliminaries
In theory, it is possible to retrieve the most semantically relevant videos from all available short videos on the internet to overcome the challenges of open-vocabulary text descriptions in the industry. However, in academia, in order to efficiently verify the feasibility of our method and to advance this research direction, we gather and filter a vast human-centric video database (HcVD) consisting of 425,988 videos, sourced from [11, 28, 29, 30, 31, 32, 33].
To train a better retrieval model on human-centric videos, we leverage the widely used Qwen2-VL [34] model to synthesize textual descriptions following Zhao et al [35]. We clarify that the synthetic captions are used solely to train the retriever, and are not involved in the motion RAG pipeline. VimoRAG is a ready-to-use framework that requires neither large-scale annotated videos nor text-to-text retrieval. Additionally, we enhance the dataset quality by employing AlphaPose [36] to filter out videos without human detection. More details are presented in the Appendix C.
@@ 3.2 Gemini Motion Video Retriever
Figure 3 illustrates the overall architecture of our method. The model builds upon the CLIP [37] framework, where the text and video branches are encoded separately, and their similarity is computed for subsequent ranking.
Our model consists of two independent retrievers—object-level and action-level—whose outputs are fused by a lightweight router to produce the final similarity score. The object-level retriever captures visual entities (objects) and their textual arguments, while the action-level retriever targets motion and predicate-level semantics.
@@
Specifically, for the action-level retriever, we carefully design an action encoder $\theta_{\mathcal{A}}$ and predicate semantic extractor $\theta_{\mathcal{P}}$. In the action encoder, we first extract the 2D human keypoints from the $m$ frames of an input video $v$. Then we encode each frame of keypoints to the feature space separately. After this, we add learnable position embeddings frame by frame and feed them into a temporal encoder. In order to avoid information loss in the sequential encoding process, the residual operation is adopted here. To obtain a good initial weight, we adopt the pretrained AlphaPose [36] and the pretrained MotionBERT [38] as the keypoints detector and encoder accordingly. The temporal encoder is implemented by a transformer block [19] and is initialized randomly. $\theta_{\mathcal{P}}$ is initialized from the text encoder in InternVideo [39].
We train the action-level retriever using the contrastive learning loss [40] $\mathcal{L}_{action}=\mathcal{L}_{p2a}+\mathcal{L}_{a2p}$ ($\mathcal{L}_{a2p}$ is provided in the Appendix D due to space limitations):
@@
where $\mathbf{p}$ and $\mathbf{a}$ are embedding vectors encoded by $\theta_{\mathcal{P}}$ and $\theta_{\mathcal{A}}$ for the input value $t$ and $v$, respectively. $B$ denotes the number of text-video pairs in a training batch. $s(,)$ denotes the cosine similarity. For the object-level retriever, we directly adopt InternVideo [39], one of the most widely used VFMs, owing to its extensive common knowledge. The argument semantic extractor $\theta_{\mathcal{G}}$ and object encoder $\theta_{\mathcal{O}}$ are initialized using the text encoder and video encoder of InternVideo. During the fine-tuning stage, we utilize a loss function $\mathcal{L}_{object}$ similar to $\mathcal{L}_{action}$. Let g and o denote the embedding vectors encoded by $\theta_{\mathcal{G}}$ and $\theta_{\mathcal{O}}$ for the input value $t$ and $v$, respectively. We just replace p with g and replace a with o in $\mathcal{L}_{action}$, and then we can get the symmetric loss function $\mathcal{L}_{object}$. It is worth emphasizing that the two semantic extractors do not explicitly extract the semantics of predicates and arguments. We hypothesize that, through contrastive learning with different video features, each semantic extractor implicitly captures its respective focus—predicates and arguments.
@@
Figure 4: The McDPO training strategy. Given a text $t$ and a retrieved video $v$, we first perform visual demonstration-enhanced instruction tuning to establish a base reference model $\pi_{ref}$. Then, based on the motion-centric dual-alignment reward model, we construct a preference dataset and apply DPO training. The reward model jointly measures motion similarity in the feature space and semantic consistency with the text, guiding the model to learn informative motion priors and maximize preference rewards through self-improvement.
@@ 3.3 Motion-centric Dual-alignment DPO Trainer
Given an input text $x=\{x_{1},x_{2},...x_{n_{x}}\}$ of $n_{x}$ tokens, a system prompt $\mathcal{S}$ of the LLM, an instruction template $\mathcal{T}$, and a retrieved video $v=\{v_{1},v_{2},...,v_{n_{v}}\}$ of $n_{v}$ frames, we first embed $v$ into $k$ segment-wise embeddings $E^{v}=\{E^{v}_{1},E^{v}_{2},...,E^{v}_{k}\}$ following Maaz et al. [10] We then concatenate these elements to obtain the final input embeddings $E^{f}=[emb(\mathcal{P}),\mathcal{T}(emb(x),E^{v})]$ where $emb(\cdot)$ denotes the embedding layers of the LLM. For the target motion, we leverage the widely used VQ-VAE [8] to encode the continuous motion sequence into discrete motion tokens $y=\{y_{1},y_{2},...,y_{n_{y}}\}$ following MotionGPT [3]. Inspired by them, we fine-tune the LLM using the following instruction-tuning format in Stage 1: System Prompt $\mathcal{P}$: You are a helpful AI assistant.
@@
MoMask [13] 5.982±.089 0.110±.003 0.195±.006 0.266±.006 5.625±.023 7.558±.119
T2M-GPT [1] 5.359±.078 0.108±.006 0.186±.005 0.255±.006 5.773±.037 7.648±.100
MDM [20] 5.907±.107 0.113±.004 0.200±.004 0.278±.004 6.013±.020 8.131±.080
MotionDiffuse [2] 5.485±.038 0.110±.002 0.194±.002 0.266±.003 6.038±.005 6.884±.095
MLD [22] 5.410±.085 0.114±.003 0.200±.005 0.270±.004 6.005±.029 7.558±.086
MotionGPT [4] 6.202±.186 0.087±.005 0.151±.007 0.209±.008 6.640±.025 7.684±.111
ReMoDiffuse [6] 9.639±.069 0.110±.004 0.188±.006 0.256±.005 5.465±.015 7.540±.120
@@
To construct such a training set, we first use $\pi_{ref}$ to randomly sample $\kappa$ times to generate a motions’ candidate set. In contrast to recent works [41, 42], which rely on human or AI-based proxies as reward models, we devise a more efficient dual-alignment reward model. This model outputs a reward score for a sampled motion sequence $\hat{y}_{i}$ as follows:
@@
where $\ell(\cdot)$ and $d(\cdot)$ denote the distribution distance and Euclidean distance computed based on the features of the two inputs, respectively. $w_{\ell}$ and $w_{d}$ are hyper-parameters to respectively control the degree of alignment within the motion modality and between the text-motion modalities. The reward model encourages the preferred motions to be closer to the reference motions in the feature space and more semantically aligned with the paired text in the semantic space. By leveraging the reward model, we collect the chosen motions $y_{w}$ and the rejected motions $y_{l}$ from the existing motions’ candidate set, thereby constructing a DPO dataset $\mathcal{D}_{dpo}=\{(x,v,y^{w},y^{l})\}$. Finally, we adopt the following DPO training objective [43, 44]. Here $\pi_{\theta}$ and $\sigma$ denote the policy model and the logistic function, respectively. $\gamma$ here denotes the weighting coefficient.
@@
We employ phi-3-mini [9] as the backbone LLM and utilize the LoRA [45] tuning method in both training stages. In Stage 1, we additionally tune the visual adapter while freezing the remaining modules. In Stage 2, all modules except for the LLM are kept frozen. Due to space constraints, further training configurations are detailed in the Appendix D.
Table 2: Results on HumanML3D test set. “*” denotes results from original papers; others are reproduced using official code. VimoRAG achieves the best FID and competitive performance across metrics among existing motion LLMs. This framework significantly improves five metrics (highlighted in red) over MotionGPT [3] (Phi3-3.8B), demonstrating the substantial advantage of incorporating video priors. The complete results with confidence intervals are shown in Table 7.
@@
MoMask [13] 0.048 0.519 0.715 0.809 2.955 9.632
T2M-GPT [1] 0.112 0.489 0.679 0.774 3.125 9.691
MDM [20] 0.454 0.419 0.606 0.712 3.636 9.449
MotionDiffuse [2] 0.672 0.492 0.685 0.784 3.085 9.499
MLD [22] 0.425 0.468 0.656 0.759 3.266 9.698
ReMoDiffuse [6] 0.125 0.493 0.676 0.775 3.047 9.211
LMM* [27] 0.040 0.525 0.719 0.811 2.943 9.814
MotionLab* [46] 0.167 0.810 2.912 9.593
MotionLCM* [47] 0.304 0.502 0.698 0.798 3.012 9.607
MotionCLR* [48] 0.269 0.544 0.732 0.831 2.806
MotionGPT* [4] 0.232 0.492 0.681 0.778 3.096 9.528
BiPO* [49] 0.030 0.523 0.714 0.809 2.880 9.556
StableMoFusion* [50] 0.098 0.553 0.748 0.841 9.748
MoGenTS* [51] 0.033 0.529 0.719 0.812 2.867 9.570
LAMP* [52] 0.032 0.557 0.751 0.843 2.759 9.571
@@
MotionGPT-2* [25] Llama3-8B 0.191 0.496 0.691 0.782 3.080 9.860
MotionLLM* [53] GPT4+Gemma-2B 0.230 0.515 0.801 2.967 9.908
Wang et al.[54] Llama2-13B 0.166 0.519 0.803 2.964
ScaMo* [55] codesize 512-3B 0.617 0.443 0.627 0.734 3.340 9.217
AvatarGPT* [56] Llama-13B 0.567 0.389 0.539 0.623 9.489
MotionGPT* [3] Llama-13B 0.567 3.775 9.006
MotionGPT [3] Phi3-3.8B 0.501 0.396 0.575 0.673 3.724 9.475